How do you solve log(2x+3)-logx=1
Answer by stanbon(48535)

How do you solve log(2x+3)-logx=1
---
log[(2x+3)/x] = 1
(2x+3)/x = 10^1
2x+3 = 10x
8x = 3
x = 3/8
=================
Cheers,
Stan H.